### v3

# SH2: Self-Highlighted Hesitation Helps You Decode More Truthfully

Abstract
Large language models (LLMs) demonstrate great performance in text generation. However, LLMs are still suffering from hallucinations. In this work, we propose an inference-time method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 is based on a simple fact rooted in information theory that for an LLM, the tokens predicted with lower probabilities are prone to be more informative than others. Our analysis shows that the tokens assigned with lower probabilities by an LLM are more likely to be closely related to factual information, such as nouns, proper nouns, and adjectives. Therefore, we propose to “highlight” the factual information by selecting the tokens with the lowest probabilities and concatenating them to the original context, thus forcing the model to repeatedly read and hesitate on these tokens before generation. During decoding, we also adopt contrastive decoding to emphasize the difference in the output probabilities brought by the hesitation. Experimental results demonstrate that our SH2, requiring no additional data or models, can effectively help LLMs elicit factual knowledge and distinguish hallucinated contexts. Significant and consistent improvements are achieved by SH2 for LLaMA-7b, LLaMA2-7b and Mistral-7b on multiple hallucination tasks.  [^1] We release our code at https://github.com/0-KaiKai-0/SH2.
SH2: Self-Highlighted Hesitation Helps You Decode More Truthfully
Jushi Kai  Hai Hu  Zhouhan Linthanks:   Corresponding author Shanghai Jiao Tong University {json.kai, hu.hai}@sjtu.edu.cnlin.zhouhan@gmail.com
1 Introduction
Depending on massive training corpora, large language models (LLMs) have made tremendous progress in natural language understanding and text generation (Touvron et al., 2023a, Jiang et al., 2023, OpenAI, 2023). However, during reasoning and generation, LLMs could suffer from hallucinations and generate non-factual answers (Zhang et al., 2023b).
Existing approaches address this problem mainly by targeting the generalization error during inference. It has been found that a problematic alignment process could mislead LLMs into hallucination (Lin et al., 2022, Zhang et al., 2023a). In the domain not included in the training, LLMs assign similar probabilities to correct and wrong choices since they do not have enough relevant knowledge to distinguish them.
To fill the gap of knowledge, retrieval augmentation methods Peng et al. (2023); Gao et al. (2023); Gou et al. (2023) leverage external knowledge bases and tools to correct the output of LLMs. Although they provide additional information for LLMs, it still seems impossible to guarantee zero error in the external knowledge. Other researchers propose decoding reformulation methods Li et al. (2023b); O’Brien and Lewis (2023); Chuang et al. (2023) to address hallucinations. However, most of them rely on external data with human labor or powerful models to rescale probability distribution during decoding.
Recent works Wei et al. (2022); Goyal et al. (2023) have demonstrated that a few more computation steps for LLMs can make a difference. Wei et al. (2022) propose chain-of-thought prompting to elicit intermediate reasoning steps of LLMs. These intermediate steps can lead to better answers. However, elaborate prompts need to be prepared in advance. Goyal et al. (2023) append learnable pause tokens to the input prefix and train models from scratch. These delays introduced by the pause tokens provide the model with more computation steps to generate better answers.
In this paper, we seek to leverage the LLM itself to highlight the informative tokens and digest them as input during the hesitation steps to elicit truthful knowledge inside LLMs. We propose a simple yet effective method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 introduces hesitations to give LLMs more time to understand contexts and answer questions. For LLMs, the tokens assigned with lower probabilities are harder to predict, while more likely to be informative. LLMs can select these key tokens by themselves from the input and hesitate on these highlighted tokens. We calculate the difference brought by highlighted tokens through contrastive decoding (Li et al., 2023c) and integrate it into the output probability. Experiments on multiple tasks demonstrate that such a difference could elicit knowledge about the facts inside the model and significantly mitigate hallucinations. Unlike other methods, our method does not leverage any other external tools or data. Additionally, it can be directly deployed during inference with no more training.
2 Related Work
Since training approaches have high requirements for computational resources, many inference-time approaches are proposed to circumvent LLM hallucinations. These methods can also be divided into retrieval augmentation and decoding reformulation.
Retrieval augmentation approaches resort to external knowledge bases and tools to help correct hallucinations. Additional information is retrieved to provide relevant knowledge for LLMs and support their generation. Peng et al. (2023) and Gao et al. (2023) leverage search engines to attribute and refine the output of language models. Furthermore, Gou et al. (2023) enable multiple tools to correct responses of LLMs autonomously during the interaction with external tools.
Decoding reformulation approaches work on reformulating the probability distribution of the outputs. Li et al. (2023b) introduce Inference-Time Intervention (ITI) to locate truthful directions of TruthfulQA Lin et al. (2022) and shift model activations toward truthfulness during inference. Despite no additional training data, they still need to train a domain-specific classifier for each attention head with the data of TruthfulQA. While Li et al. (2023c) propose Contrastive Decoding (CD) to capture the likelihood difference between large and small models. They suggest that this difference signals which texts should be preferred when generating. And O’Brien and Lewis (2023) leverage contrastive decoding to improve reasoning quality for LLMs. However, the original contrastive decoding must utilize a strong LLM as the expert model. Chuang et al. (2023) propose DoLa to localize the premature layer in LLaMA Touvron et al. (2023a), which contains the least knowledge about the input text. They use the last layer as the expert model and the premature layer as the amateur model, and contrast the prediction probability between them. Shi et al. (2023) leverage contrastive decoding to amplify the output difference between with and without contexts for summarization tasks.
The idea of using extra decoding steps while predicting harder tokens can be dated back to Adaptive Computation Time proposed by Graves (2017). Shin et al. (2020) introduce AUTOPROMPT to combine the original inputs with trigger tokens to elicit knowledge from pretrained models. Wiegreffe et al. (2022) investigate that additional steps to generate rationales could lead to a more faithful model. More recently, Goyal et al. (2023) have also demonstrated that for LLMs, inserting extra computation steps to allow the model to pause before generation can enhance the performance on question answering and reasoning tasks. Nevertheless, their method only works when the model is both pre-trained and finetuned with pauses.
Figure 1: The pipeline to construct and leverage our Self-Highlighted Hesitation.
3 Self-Highlighted Hesitation
The illustration of our Self-Highlighted Hesitation is shown in Figure 1. For the original inference procedure, the instruction, document, and summary are directly fed into the LLM to generate the judgment. The LLM can be easily confused by the hallucinated context.
For our SH2, we underline the key tokens of the input document by the prediction likelihood. These tokens are listed as a hesitation and appended to the document. The prediction probability is scaled with the difference of confidence between the hesitation-prompted input and the original input. Our method can effectively help LLMs identify the hallucinated context. We will elaborate our method in the following subsections.
3.1 Key Tokens
We select key tokens based on the prediction probability given by the LLM. The decoding procedure of a language model $\theta$ can be formalized as:
$\hat{p}(x_{t})=p_{\theta}(x_{t}|x_{<t})$ (1)
where $x$ is the context, $t$ denotes the current predicting position of $x$ and $p_{\theta}$ gives the prediction probability of the token $x_{t}$ by the model $\theta$. We obtain the probability of generating $x_{t}$ by feeding previous tokens to the model.
The probability measures the confidence of the language model for each token given the previous context. It represents how simple it is to infer the token from the previous context by the model. The tokens with the lowest probabilities bring the most semantic information and are the hardest to predict. We regard these tokens as Key Tokens. They deserve more attention from the language model and help comprehend the whole context.
Figure 2: The heatmap to show the normalized top-$\eta$ recall for the top 20 most frequent POS tags. The light color and the high value indicates that these POS tags occupy high proportions in the hardest part. 1000 documents are sampled from the summarization track of HaluEval Li et al. (2023a), which is a dataset collected from CNN/Daily Mail Hermann et al. (2015). We extract the hardest words that contain key tokens from these documents with the proportion of $\eta$ ranging from 1% to 10% by LLaMA2-7b Touvron et al. (2023b).
3.2 Relation between Key Tokens and Factual Information
To illustrate that key tokens selected in this fashion are closely related to the factual knowledge, we analyze the Normalized Top-$\eta$ Recall for POS (part-of-speech) tags from the perspective of grammar. It measures the percentage of POS tags in the document that appear in the hardest part.
For the document $X_{i}$, the number of words with POS tag $z_{k}$ is:
$N(X_{i},z_{k})=\#\{POS(X_{i})=z_{k}\}$ (2)
where $POS(\cdot)$ is the function to derive POS tags and #$\#\{POS(X_{i})=z_{k}\}$ is to count how many words in $X_{i}$ has the POS tag of $z_{k}$.
For a dataset with $m$ documents, the recall of POS tag $z_{k}$ can be calculated by:
$\Delta_{\eta}(z_{k})=\frac{\sum_{1\leq i\leq m}N(T(X_{i},\eta),z_{k})}{\eta\cdot\sum_{1\leq i\leq m}N(X_{i},z_{k})}$ (3)
where $T(X_{i},\eta)$ is the set of words that are among the lowest $\eta$ portion of probability predicted by the language model in $X_{i}$. The $\eta$ in the denominator is to normalize the scale of $\Delta_{\eta}(z_{k})$
$\Delta_{\eta}(z_{k})$ measures how hard the words are with the POS tag $z_{k}$ for LLMs to predict with respect to the proportion $\eta$. $\Delta_{\eta}(z_{k})$ is larger than 1 means that $z_{k}$ is concentrated in the hardest part of documents. Specifically, $\Delta$ will be 1 for all the tags when $\eta$ is set to be 100%. Figure 2 demonstrates the normalized top-$\eta$ recall of the most frequent POS tags for the summarization track of HaluEval Li et al. (2023a).
It can be observed from Figure 2 that content words such as adjectives (JJ), nouns (NN), proper nouns (NNP), adverbs (RB) and conjugated verbs (VBD, VBG, VBP) are more frequent in the hardest words we extract, as the light color in the heatmap indicates. On the other hand, function words such as conjunctions (CC), determiners (DT), and prepositions (IN) are less frequent, as indicated by the darker colors. [^2] The base form of the verbs (VB) seems to be an exception. It is probably because VBs have less information load than conjugated verb forms, as the latter also encode tense, aspect, and other information about the verb. This suggests that the normalized top-$\eta$ recall is positively correlated with factual informativeness. Therefore, picking out key tokens that carry more content, rather than those that serve as grammatical functions, could intuitively help the model to focus on factual information. Our method can thus be understood as a way to “highlight” the content words to LLMs, potentially making them less likely to hallucinate, the effect of which will be evaluated in our experiments. More cases will be given in Appendix D.2.
3.3 Construction of Hesitations
To this end, we attempt to solve hallucinations by highlighting the key tokens that are hard to predict. For the text sequence $X=(x_{0},x_{1},\dots,x_{n})$, we can obtain the corresponding probability sequence by the language model:
$\hat{P}(X)=(\hat{p}(x_{1}),\hat{p}(x_{2}),\dots,\hat{p}(x_{n}))$ (4)
We use the following strategy to select key tokens according to their probability $\hat{P}(X)$ and construct hesitations:
First, construct a candidate key-token set by selecting $\eta\cdot n$ tokens with the lowest prediction probabilities of $\hat{P}(X)$, where $\eta\in(0,1)$ is the sampling proportion we preset for $X$. Then, derive the target key-token set $T(X)$ by randomly drop out $\lambda\cdot\eta\cdot n$ tokens from the candidate set, where $\lambda\in[0,1]$ is the drop-out rate.
For long documents, to avoid $T(X)$ got dominated by one or two POS tags, we introduce the drop-out rate $\lambda$ to randomly select key tokens among a larger pool of candidates, whose size is determined by $\eta$. As a result, $(1-\lambda)\cdot\eta\cdot n$ tokens are retained in $T(X)$ to fill as input during the hesitation steps.
Finally, construct the Hesitation text $H(X)$ with the target key-token set $T(X)$ for the input context: $H(X)=``Pondering:<T(X)>."$, where ``$Pondering:"$ is the prefix. We keep the tokens’ order of $T(X)$ in which they appear in the original text $X$.
With the hesitation text following the original text, the language model can focus more on key tokens and have more time to infer the non-hallucinated answer.
3.4 Contrastive Decoding on Hesitations
Using hesitations as prompting or data augmentation may not be enough. Since we do not introduce new factual information in the hesitation text, the improvement from the hesitation-prompted input might be limited. Therefore, we resort to what difference hesitations can bring during decoding. Different from previous work Li et al. (2023c); Chuang et al. (2023); O’Brien and Lewis (2023) that contrasts between models, we use contrastive decoding (Li et al., 2023c) from the perspective of data:
$p_{\mathbf{CD}}(Y|X)=\text{softmax}(\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)})$ (5)
where $Y$ is the output text and $X^{\prime}$ is the concatenation of the original input text $X$ and its hesitation $H(X)$.
The difference of confidence is traded off with the base probability $p_{\theta}(Y|X^{\prime})$:
$p_{\mathbf{H}}(Y|X)=\left\{\begin{array}[]{cl}\text{softmax}(p_{\theta}(Y|X^{\prime})\cdot(\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)})^{\alpha}),\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\alpha\neq 0\\ p_{\theta}(Y|X^{\prime}),\text{ otherwise }\end{array}\right.$ (6)
where $\alpha$ is a hyper-parameter used for scaling the difference of confidence between with and without hesitations. When $\alpha=0$, $p_{\mathbf{H}}(Y|X)$ is equal to the base probability $p_{\theta}(Y|X^{\prime})$, which means the model directly decodes with the input text and the hesitation.
With $p_{\mathbf{H}}(Y|X)$, the confidence difference could be noticed, especially when the prediction probability is low. While for the high-confidence tokens, the probability change brought by hesitations might be marginal. $p_{\mathbf{CD}}(Y|X)$ could be very uniform. Therefore, LLMs can mainly follow the base probability $p_{\theta}(Y|X^{\prime})$ to make predictions.
4 Experiments
We conducted experiments on five tracks of three hallucination evaluation benchmarks.
Models TruthfulQA FACTOR HaluEval-Sum
MC1 MC2 MC3 Truth Truth*Info Wiki News Acc-A Acc-H
LLaMA-7b 23.62 41.21 19.33 30.97 27.78 58.55 58.40 26.06 18.94
  +Alpaca 26.93 42.97 19.79 39.17 38.92 57.11 58.20 37.24 18.31
  +ITI* 25.9 - - 49.1 43.5 - - - -
  +13b-CD* 24.4 41.0 19.0 55.3 44.4 64.4 62.3 - -
  +DoLa 31.95 52.21 28.17 40.88 39.66 61.96 61.68 - -
  +SH2 (Ours) 27.91 55.63 29.73 64.99 41.49 63.06 65.54 31.36 26.80
LLaMA2-7b 28.40 43.39 20.53 48.59 40.76 58.65 72.20 48.03 19.88
  +SH2 (Ours) 33.90 57.07 29.79 64.38 42.23 64.09 73.65 50.56 50.41
Mistral-7b 31.58 48.14 23.89 - - 60.72 75.97 41.03 40.79
  +SH2 (Ours) 30.84 52.52 27.39 - - 60.86 77.03 42.87 42.36
Table 1: Truthfulness scores (%) on the three benchmarks. We also underline the second-best scores SH2 achieved. For ITI (Li et al., 2023b), "*" means we report the results on TruthfulQA from their paper since they use 5% of TruthfulQA questions in few-shot prompting. For 13b-CD, "*" means we report the results of Contrastive Decoding (CD) (Li et al., 2023c) from Chuang et al., 2023. They use LLaMA-7b as the amateur model and LLaMA-13b as the expert model. We maintain the same experimental settings for the other baselines and our SH2 in our implementation.
4.1 Benchmarks
TruthfulQA
TruthfulQA (Lin et al., 2022) is a benchmark to measure the truthfulness of a language model in question answering. It has 817 samples and both tracks of generation and discrimination for evaluation. To automatically evaluate the generation quality of LLMs, it introduces GPT-judge, a fine-tuned GPT-3. We use “Truth” to represent the percentage of the generated answers that are true and “Truth*Info” for answers that are both true and informative.
For the discrimination track, it offers sets of true and false reference answers for each question. We compute the likelihood of each answer given the question, and compare the probabilities of true answers against false answers to derive MC1, MC2 and MC3 scores. The definition of the MC (Multiple-Choice) metrics can be referred to in Appendix B.
FACTOR
FACTOR (Muhlgay et al., 2023) puts more attention on the consistency of contexts and measures the tendency of a language model to generate factual information. It is a text completion task to identify the correct completion from the non-factual statements given the prefix. It contains two datasets of different sources: Wiki-FACTOR and News-Factor. There are 2994 and 1036 examples in each dataset. We gauge the factuality by whether the language model assigns the highest likelihood to the factually correct completion over the other options.
HaluEval-Sum
HaluEval (Li et al., 2023a) provides texts, each paired with a hallucinated and right responses. We use its summarization track to evaluate LLMs’ truthfulness on longer sequences. It has 10000 samples. For each sample, we ask LLMs to judge whether the provided summary contains non-factual or hallucinated information against the given document. We compute accuracy for hallucinated summaries and right summaries respectively. Arithmetic-mean accuracy (Acc-A) and harmonic-mean accuracy (Acc-H) are reported in our experiments.
4.2 Experimental Settings
We apply our SH2 on LLaMA-7b (Touvron et al., 2023a), LLaMA2-7b (Touvron et al., 2023b) and Mistral-7b Jiang et al. (2023). We compare it with other SOTA (state-of-the-art) methods, including Alpaca (Taori et al., 2023), ITI (Li et al., 2023b), 13b-CD (Chuang et al., 2023) and DoLa (Chuang et al., 2023). All of these baselines use LLaMA-7b as their backbone. It should be noted that ITI uses 5% samples of TruthfulQA in few-shot prompting. Additionally, LLaMA-13b is used in 13b-CD to be contrasted with the 7b model. We implement Alpaca and DoLa following the official instructions and report the evaluation results of our implementation.
We append hesitations to the original inputs for TruthfulQA and HaluEval-Sum. As for FACTOR, hesitations are prepended to inputs. Because FACTOR is a task of completing articles, it hurts the continuity of articles to insert hesitations.
Since there are only about a dozen tokens in each question of TruthfulQA, the sample proportion $\eta$ of LLaMA-7b is set to be 10% and 40% for the discrimination track and the generation track respectively. And the drop-out rate $\lambda$ is set to be 0. While for HaluEval-Sum, which has about a thousand tokens in each document, $\eta$ and $\lambda$ are set to be 6% and 0.33 for LLaMA-7b. The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Appendix A.
4.3 Main Results
The results on the three benchmarks are shown in Table 1. Our SH2 obtained significant and consistent improvements for LLaMA-7b, LLaMA2-7b and Mistral-7b. When compared with other SFT or decoding reformulation methods, our method outperforms them in most of the metrics of the benchmarks. There is no additional data or model used in our SH2. We ask models to select the hardest tokens and hesitate on them. Although LLaMA2 and Mistral have been aligned to generate truthful answers during its training procedure, they still benefit greatly from our inference-time method.
Moreover, our SH2 achieved SOTA on both the generation and the discrimination tracks of TruthfulQA. The scores of our method are also the highest or the second best on the other three discrimination tasks. It can be concluded from the table that our approach can effectively elicit factual knowledge inside LLMs. It can not only help LLMs distinguish factual contexts and hallucinated contexts, but also guide them to generate more truthful answers.
Models Precision Recall F1
LLaMA-7b 17.10 12.44 14.40
  +SH2 34.95 43.31 38.69
LLaMA2-7b 42.55 11.26 17.81
  +SH2 50.59 47.78 49.15
Mistral-7b 41.56 44.2 42.84
  +SH2 43.48 47.54 45.42
Table 2: Precision, recall and F1 scores (%) on HaluEval-Sum.
4.4 Bias in HaluEval-Sum
Looking into the summarization track of HaluEval Li et al. (2023a), it can be noted from Table 1 that there is a huge difference between the values of Acc-A and Acc-H for LLaMA-7b, Alpaca and LLaMA2-7b. Acc-A is the average of the accuracies on hallucinated summaries and right summaries. Acc-H is computed by averaging the reciprocals of the accuracies and then taking the reciprocal of the average. Acc-A can be affected by extreme values easily, while Acc-H is more balanced.
The great gap between Acc-A and Acc-H denotes LLMs’ bias in hallucinated summaries and right summaries even for LLaMA2-7b. While our method effectively alleviates the problem.
Furthermore, we regard hallucinated summaries as positive labels and right summaries as negative. We also calculate precision, recall and F1 scores to evaluate the truthfulness more thoroughly. The scores are reported in Table 2. Precision denotes the percentage of real hallucinations the model determines to be hallucinated, while recall denotes the accuracy on hallucinated summaries. It can be learned from the table that LLaMA2-7b has more confidence in identifying factual summaries. However, it is still challenging for LLaMA2-7b not to be confused by hallucinated summaries, which is a potential risk for the development of LLMs.
Nevertheless, our method successfully mitigates this imbalance and benefits the general performance. It can advance the discriminant ability of models comprehensively.
(a) Average scores (%) for different highlighted tokens with the effective sampling proportion $\eta^{\prime}$ ranging from 1% to 8%. The errorbar denotes standard deviations.
(b) Harmonic accuracy (%) for different highlighted tokens with respect to the effective sampling proportion $\eta^{\prime}$.
Figure 3: Different choices of highlighted tokens by LLaMA2-7b.
5 Analysis
5.1 Choices of Highlighted tokens
Key tokens are sampled and highlighted by how hard they are for large language models to predict. In order to verify the effect of key tokens, we compare the performance of SH2 with different choices of highlighted tokens.
In contrast to key tokens that are the hardest to predict, we sample the easiest tokens with the highest prediction probability. Additionally, we also randomly sample the same number of tokens for comparison. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The effective sampling proportion $\eta^{\prime}=(1-\lambda)\cdot\eta$ ranges from 1% to 8% with the step of %1. Average scores and standard deviations are calculated for the three choices of highlighted tokens.
The accuracies on hallucinated summaries and right summaries, and overall scores (Acc-A and Acc-H) are shown in Figure 3a. When the hardest tokens are highlighted in hesitations, the LLaMA2-7b obtains better and more consistent performance than when highlighting the easiest tokens or randomly. It can be noticed that the accuracy on the hallucinated summaries is much better than accuracies of the other two.
Figure 3b illustrates the harmonic accuracy for different choices of highlighted tokens with respect to the sampling proportion $\eta$. The scores of the hardest tokens are consistently better than the other two. It is worth noting that only 1% key tokens to be highlighted in hesitations are helpful enough for language models to distinguish hallucinated contexts.
Figure 4: Precision, recall and F1 scores (%) for different highlighted tokens with the effective sampling proportion $\eta^{\prime}$ ranging from 1% to 8%. The errorbar denotes standard deviations. The scores of the vanilla LLaMA2-7b are obtained by evaluating on the whole dataset of HaluEval-Sum.
Moreover, we compare precision, recall and F1 scores for the three choices of highlighted tokens in Figure 4. The disturbance brought by highlighting the easiest tokens or random tokens hurts the precision of LLMs, while the hardest tokens are beneficial. It can also be concluded from the figure that the manner of hesitations helps mitigate the bias in HaluEval-Sum.
5.2 Manners of Hesitations
In addition to underlining key tokens in the input text, we can also repeat the text or pause. Experiments are conducted on TruthfulQA with LLaMA-7b. For the repetition manner, we can repeat the questions in hesitations. While for the pausing manner, key tokens $T_{2}(X)$ are replaced by several pause words (".") in hesitations. The performance of pausing hesitations is evaluated with 3, 6, 9 and 12 pause words. MC2 scores with different numbers of pause words and the adjustment hyper-parameter $\alpha$ are shown in Figure 6. 6 pause words are used in hesitations in the following study.
The scores on the discrimination track and the generation track of the three manners are reported in Table 3. Hesitations with key tokens achieve the most significant improvement on both tracks, while pausing hesitations even hurt the generation quality of LLaMA-7b. It can be found from the table that all of the three manners help discriminate hallucinated answers. The results demonstrate that the difference brought by hesitations can elicit factual knowledge inside language models.
Models MC1 MC2 MC3 Truth*Info
LLaMA-7b 23.62 41.21 19.33 27.78
  +key tokens 27.91 55.63 29.73 41.49
  +pauses 27.54 48.64 24.95 22.52
  +repetition 26.93 45.05 21.16 31.70
Table 3: Multiple-choice and generation scores (%) on TruthfulQA for different manners of hesitations.
5.3 Effect of Contrastive Decoding
Figure 5: Effect of contrastive decoding on hesitations. The dashed line in gray represents the MC2 score of the vanilla LLaMA-7b. Dashed lines in other colors represent the MC2 scores for the standard contrastive decoding with different numbers of highlighted tokens.
Our method does not use the original form of contrastive decoding. The difference of confidence brought by hesitations, which is the item of contrastive decoding, is traded off with the base probability. We conduct experiments on TruthfulQA with different values of the adjustment hyper-parameter $\alpha$ to study the effect of contrastive decoding (CD).
The effect of contrastive decoding with different numbers of highlighted tokens is shown in Figure 5. Since there are only about a dozen tokens in each question of TruthfulQA, the hesitation contains only one token when sampling the top 10% hardest tokens. LLaMA-7b benefits a lot from the one or two highlighted tokens.
When $\alpha$ is set to be 0, the model directly decodes with the input text followed by the hesitation. The MC2 score even drops a little compared to the vanilla LLaMA-7b. As more weight is put on the difference brought by hesitations, the model obtains a more significant improvement. Specifically, when the weight is big enough to ignore the base probability, the decoding procedure is equivalent to standard contrastive decoding without the base probability as the equation (5). It can be concluded from the figure that, the highlighted tokens make a difference in the confidence to identify factual knowledge.
6 Conclusion
In this paper, we study the bias in capturing important factual information for large language models and introduce a novel inference-time method SH2. Our method proposes to highlight the key tokens that are hard for LLMs to predict and construct hesitations with these informative tokens. To help LLMs capture critical semantic information, the decoding procedure is reformulated with the probability difference brought by hesitations. Comprehensive experiments and analysis are conducted on multiple tasks. SH2 significantly improves the truthfulness of LLMs with no need for external data or models. In the age of LLMs, our SH2 could inspire a new paradigm for faithful LLMs.
Limitations
Existing research on LLMs’ hallucination does not pay much attention to other dimensions of the generation quality. Despite experiments on the generation track of TruthfulQA, we have yet to explore the diversity and soundness of the generated contents by our SH2, which is closely related to the generalization ability. It is worth studying how to optimize LLM’s truthfulness and generalization simultaneously.
Besides, our method is an inference-time method without leveraging external data or models. Consequently, it could be promising to integrate our method with other data-enhanced or model-enhanced methods. The idea of constructing hesitations can also be applied in retrieval augmentation approaches.
Ethics Statements
Our work pertains to large language models’ hallucinations. In this work, we use only publicly available data and artifacts. There are no ethical issues in our paper, including its motivation and experiments.
Appendix A Hyper-parameter Settings
The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Table 4. The drop-out rate $\lambda$ is set to be 0 for the discrimination track and the generation track of TruthfulQA, whose questions are of a dozen tokens. It is set to be 0.33 for FACTOR and HaluEval-Sum, because these benchmarks have longer documents.
Appendix B Multiple-Choice Metrics
For the discrimination track of TruthfulQA, we use MC1, MC2 and MC3 scores to measure the truthfulness of a language model. The definitions of each score are as follows.
  • MC1: Among the set of true and false reference answers, we need to choose the best correct answer. MC1 is computed by whether the language model assigns the highest likelihood to the best correct answer over the false answers given the question.
  • MC2: MC2 is the total normalized probability of the true reference answers. The score is the probability mass for correct answers.
  • MC3: MC3 is computed by whether the language model assigns a higher likelihood to the correct answers over the false answers given the question.
Appendix C Constrastive Decoding
C.1 Effect of CD with Pause Words
Figure 6: Effect of contrastive decoding on hesitations. The dashed line in gray represents the MC2 score of the vanilla LLaMA-7b. Dashed lines in other colors represent the MC2 scores for the standard contrastive decoding with different numbers of pause words.
As shown in Figure 6, we also studied the effect of contrastive decoding with different numbers of pause words. A similar conclusion can be observed from the figure.
Figure 7: An illustration to show the effect of contrastive decoding. The options (a) $\sim$ (d) are correct answers and (e) $\sim$ (h) are false answers.
C.2 Case Study
A case study is given in Figure 7 to illustrate the effect of our self-highlighted hesitations and contrastive decoding. We calculate the probability given by LLaMA-7b. $p_{\theta}(Y|X)$ and $p_{\theta}(Y|X^{\prime})$ stand for standard decoding without and with the hesitation for each option. The difference in probability are derived by $\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)}$.
It can be learned from the figure that, with hesitations, LLaMA-7b is prone to assign higher probabilities to correct answers and lower probabilities to false answers. The model is more confident in selecting correct answers and rejecting false answers. It demonstrates that CD helps LLMs distinguish correct answers from false ones. It plays a role in separating the positive and negative answers.
Appendix D Key Tokens
D.1 Normalized Top-$\eta$ Recall
We also calculated the normalized top-$\eta$ recall between the hardest part and the whole document for the top 20 most frequent POS tags by LLaMA-7b and Mistral-7b through the equation (3). The heatmaps are shown in Figure 8. The same statistical laws can be concluded as those of LLaMA2-7b in Section 3.2.
(a) The normalized top-$\eta$ recall of LLaMA-7b.
(b) The normalized top-$\eta$ recall of Mistral-7b.
Figure 8: The normalized top-$\eta$ recall for different POS tags by LLaMA-7b and Mistral-7b.
D.2 Visualization of Token Probabilities
(a) Visualization of token probabilities estimated by LLaMA-7b.
(b) Visualization of token probabilities estimated by LLaMA2-7b.
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 9: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 1. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
(a) Visualization of token probabilities estimated by LLaMA-7b.
(b) Visualization of token probabilities estimated by LLaMA2-7b.
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 10: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 2. The color coding in the figure represents the token generation probabilities. A spectrum from red to green is utilized to represent a range from lower to higher generation probabilities, respectively.
(a) Visualization of token probabilities estimated by LLaMA-7b.
(b) Visualization of token probabilities estimated by LLaMA2-7b.
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 11: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 3. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
Figure 9 to Figure 11 illustrate three text examples where the background color of each token denotes its generation probability given by LLaMA-7b, LLaMA2-7b, and Mistral-7b. A spectrum from red to green is utilized to represent probabilities ranging from low to high.
These examples reveal a tendency for lower probability tokens to consist of content words predominantly. However, instances exist where determiners and prepositions also exhibit lower probabilities, as exemplified in Figure 9a with the words “all” and “on” in line 5, and the word “in” in line 14. This can be attributed to the interchangeable use of different words to convey identical semantic content at the beginning of a sentence or sub-clause.
Besides, tokens with higher probabilities are typically from function words. However, exceptions are noted with certain content words demonstrating high probabilities due to the extensive knowledge memorized by LLMs, such as “typhus” and “15” in the first line of Figure 11a. Additionally, some non-initial tokens, such as “nesday” and “CC” in lines 1 and 3 of Figure 10a, exhibit high generation probabilities. It reflects the determinative role of initial tokens in setting the context for subsequent non-initial tokens.
Comparing the prediction probabilities of LLaMA-7b, LLaMA2-7b, and Mistral-7b, we did not observe significant difference. Although LLaMA2-7b and Mistral-7b demonstrate superior performance in various benchmarks compared to LLaMA-7b, there is still considerable room for improvement in their grasp of factual knowledge. Therefore, it is beneficial to incorporate hesitations to make the model pay more attention to these informative key tokens.